# DDGS-CT: Direction-Disentangled Gaussian Splatting for Realistic Volume Rendering

### v1

@@ 1 Introduction
Digitally reconstructed radiographs (DRRs) are simulated (in silico) 2D X-ray images rendered from 3D computational tomography (CT) volumes. While DRRs are widely utilized in preoperative settings, such as optimizing dose delivery and in radiation oncology Lohr et al. (1997); Yang et al. (2000); Bollet et al. (2003), their potential intraoperative applications remain underexplored due to computational bottlenecks and naive modeling capability Adler Jr et al. (1999); Peters (2001). For instance, real-time multimodal registration for image-guided procedures is currently impractical because of the time-consuming process of generating DRRs and integrating them into slice-to-volume registration Van Der Bom et al. (2011).
The advent of GPU-accelerated computing has significantly improved the efficiency of radiography simulators Badal and Badano (2009); Vidal et al. (2009); Biguri et al. (2016), though they still fall short of meeting the stringent requirements of real-time applications. Recent developments in inverse graphics have led to the creation of differentiable DRR renderers (e.g., DiffDRR Gopalakrishnan and Golland (2022), X-Gaussian Cai et al. (2024), GaSpCT Nikolakakis et al. (2024)), which show promise for inverse problems such as 2D/3D CT image registration Gopalakrishnan et al. (2023). However, these methods do not fully capture the complexities of X-ray image formation. They typically use computationally efficient ray-tracing techniques that model the attenuated photon fluence at each detector pixel by accumulating Hounsfield unit (HU) values along the 3D view-ray through the voxelized CT volume. This approach, while fast, cannot account for complex noise-inducing physical effects such as Compton scattering and beam hardening Hasegawa (1990), crucial for accurate X-ray imaging.
Physics-based Monte-Carlo simulations Agostinelli et al. (2003); Jan et al. (2004); Badal and Badano (2009); Bert et al. (2013) do not suffer from these limitations. Relying on the decomposition of CT volumes into realistic materials via HU thresholding, they accurately simulate single-photon transport and probabilistically evaluate photon-matter interactions (photoelectric effect, Compton scattering, Rayleigh scattering, etc.) to determine the final attenuation and hit-region of each photon. This approach provides a more accurate representation of X-ray image formation; but the computational intensity of these methods—requiring >$10^{8}$ iterations to render an image—makes them impractical for real-time applications and inverse problems.
@@
Our approach leverages 3D Gaussian splatting (3DGS) Kerbl et al. (2023) for efficient DRR generation from CT volumes, similar to X-Gaussian Cai et al. (2024) and GaSpCT Nikolakakis et al. (2024). However, we introduce a crucial adaptation: our direction-disentangled 3DGS (DDGS) method accounts for the dependencies of X-ray image formation on light directionality. Specifically, we decompose the radiosity contribution of 3D regions into two independent components: a non-directional component modeling the isotropic X-ray interactions with the corresponding material, and a direction-dependent component approximating higher-dimensional anisotropic interactions, such as scattering. Unlike the more generic view-dependent spherical-harmonics decomposition adopted by common 3DGS solutions Fridovich-Keil et al. (2022), ours is tuned to the specific behaviors of high-energy photons. This formulation allows our method to implicitly learn and reproduce the effects of scattering on X-ray images without the need for complex photon-transport simulations at runtime.
@@ 2 Related Work
The generation of DRRs and the simulation of X-ray images have been subjects of extensive research due to their critical roles in medical imaging, particularly in radiotherapy planning and diagnostic imaging Lohr et al. (1997); Yang et al. (2000); Bollet et al. (2003); Zhang et al. (2022). In this section, we provide some insight on X-ray imaging, traditional simulation tools, and the recent integration of machine-learning techniques.
@@
X-ray imaging relies on the interactions of high-energy photons—emitted by the X-ray source of the scanner—with the matter (e.g., body part) placed in between the X-ray source of the scanner and its radiation-sensitive detector cells (dexels). Imaging contrast arises from differential absorption and scattering of X-ray photons by tissues, primarily through the photoelectric effect, Compton scattering, and, to a lesser extent, beam hardening for polychromatic scanners. The photoelectric effect involves the absorption of an X-ray photon by an atom and is more pronounced in higher atomic number materials like bone, causing them to appear whiter on X-ray images [^1] X-ray film is historically white, turning black when exposed to high-energy photon.; whereas Compton scattering occurs when an X-ray photon collides with an outer electron, resulting in the photon being deflected and losing energy, which contributes to image noise and reduced contrast Hasegawa (1990); Seibert and Boone (2005).
A variety of Monte Carlo (MC) simulation suites Agostinelli et al. (2003); Jan et al. (2004); Badal and Badano (2009); Bert et al. (2013) model these X-ray interactions and light transport. They trace the paths of a large number of individual photons as they undergo probabilistic interactions within the material. By simulating numerous photon trajectories, Monte Carlo methods account for the random nature of photoelectric absorption and Compton scattering, as well as secondary effects such as fluorescence and electron transport. These MC solutions produce highly realistic simulations of radiographic images, though at a high computational cost which makes them inadequate for most DRR use-cases. Furthermore, for these tools to render realistic DRRs from CT volumes, one must first pre-process the CT data to assign proper materials and their corresponding physical properties to each voxel, which is essential for realistic simulation. This step is often approximated by decomposing the volume into a small set of predefined materials according to user-provided HU ranges (e.g., “air” material assigned to voxels with HU value <800<-800< - 800, “lung” assigned to voxels in range ]800,200]]-800,-200]] - 800 , - 200 ], “fat” for ]200,100]]-200,-100]] - 200 , - 100 ], etc.). This process can be time-consuming and prone to errors, affecting the fidelity of MC-based DRR generation Dorgham et al. (2020).
@@
Analytical DRR rendering methods Vidal et al. (2009); Vidal and Villard (2016); Biguri et al. (2016) have been developed in parallel to the aforementioned MC suites for real-time or near-real-time applications. These methods have a much lower computational footprint, as they do not simulate individual photon interactions but instead use mathematical models to directly calculate the primary paths and approximated attenuations of X-rays. The seminal work of Siddon (1985) on computing line integrals through a discretized CT volume is still at the core of most recent DRR solutions, e.g., which proposed differentiable formulation of the rendering steps Gopalakrishnan and Golland (2022) or integrated neural networks to add realistic noise to the output Unberath et al. (2018). However, the inherent approximations in these methods can affect their accuracy, particularly in modeling artifacts present in heterogeneous tissues or complex anatomical structures, e.g., caused by scattering (non-primary light contributions).
Recently, researchers Cai et al. (2024); Nikolakakis et al. (2024) have tried to apply 3D Gaussian splatting (3DGS) techniques Kerbl et al. (2023) to DRR rendering, i.e., optimizing a cloud of 3D Gaussians to approximate voxel data, thereby enabling faster rendering with minimal accuracy loss. However, these early solutions overlooked the specific nuances of X-ray imaging, directly applying 3DGS methods designed for natural imaging Nikolakakis et al. (2024) or oversimplifying the physical properties of X-ray attenuation Nikolakakis et al. (2024), leading to suboptimal modeling.
@@ 3.1 Preliminaries
3DGS Kerbl et al. (2023) is a point-based rendering technique using 3D Gaussians to explicitly represent scenes in a more compact manner than volumetric representations. Each Gaussian $g_{i}=\{\bm{\mu}_{i},\bm{\Sigma}_{i},\bm{\alpha}_{i},\bm{f}_{i}\}$ is defined by its mean position $\bm{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ (usually decomposed into separate scale and rotation parameters), opacity $\bm{\alpha}_{i}\in\mathbb{R}$, and radiance properties $\bm{f}_{i}\in\mathbb{R}^{k}$, where $k$ depends on the light contribution model adopted. E.g., $k=(L+1)^{2}\times 3$ for anisotropic, view-dependent RGB radiance decomposed into spherical harmonics (SH) of degree $L$, a model commonly used for natural-imaging applications. Images are rendered by casting view-rays through each pixel $p$ into the scene and alpha-blending the Gaussian contributions to the final ray color $C$, as:
@@
where $N$ is the number of points sampled along the view-ray, $c_{j}$ is the radiance (computed from $\bm{f}_{j}$, e.g., via SH) of the $j$th Gaussian on the ray, and $\widehat{\bm{\mu}}_{j}$ and $\widehat{\bm{\Sigma}}_{j}$ are the image-plane projections of $\bm{\mu}_{j}$ and $\bm{\Sigma}_{j}$.
Since the aforementioned rasterization process is differentiable w.r.t. to the Gaussian parameters, the 3DGS representation of a scene can be learned via gradient-descent, given a set of 2D scene observations $\bm{I}_{\text{GT}}$ and their corresponding camera parameters. A combination of $\ell_{1}$ and SSIM Wang et al. (2004) loss functions are commonly-adopted as criterion for this iterative optimization process Kerbl et al. (2023):
@@
with $\bm{I}$ the rendered images (based on provided camera parameters), and $\lambda$ a loss-weighting hyper-parameter. The initial state is typically obtained by sampling relevant 3D points in the scene domain for the Gaussians, e.g., from a structured-from-motion (SfM) point-cloud Ullman (1979); Schonberger and Frahm (2016).
@@
The volumetric rendering performed by 3DGS methods is conceptually similar to the one performed by analytical DRR methods Siddon (1985); Gopalakrishnan and Golland (2022), i.e., aggregating the attenuation values of light-rays cast through the voxelized CT volume. This similarity has recently motivated researchers Cai et al. (2024); Nikolakakis et al. (2024) to apply 3DGS models to DRR applications, driven by the compactness of 3DGS representations compared to voxel data (addressing the $\mathcal{O}(n^{3})$ complexity inherent to voxel rendering). By optimizing a 3DGS model to approximate the visual properties of the CT volume, the resulting representation can render accurate DRRs much faster than traditional methods, making it more suitable for real-time intra-operative visualization. Consequently, the use of traditional, slower rendering solutions can be limited to generating DRR targets ($\bm{I}_{\text{GT}}$) for the offline (i.e., pre-operative) training.
However, these early 3DGS-for-DRR solutions Siddon (1985); Gopalakrishnan and Golland (2022) do not account for noise-inducing photon interactions (e.g., scattering) when applying the analytical methods Sharp et al. (2010); Biguri et al. (2016) to render their training data. Since scatter-free DRRs are not affected by ray directions (i.e., the attenuation at each point is naively considered isotropic), Cai et al. (2024) simplified the radio-intensity function of each Gaussian to $c_{i}=\operatorname{sigmoid}{(\bm{b}\cdot\bm{f}_{i})}$, where $\bm{b}\in\mathbb{R}^{k}$ is a direction-independent optimizable basis vector shared by all splats. Furthermore, their Gaussian point-cloud is initialized via evenly-spaced Cai et al. (2024) or random-uniform Nikolakakis et al. (2024) sampling of the scanned 3D space, i.e., ignoring relevant geometrical and material properties of the target CT data. Inadequate initialization strategies negatively impact the convergence of 3DGS models and the resulting image accuracy Jung et al. (2024). In this work, we propose to reformulate both the initialization strategy and the radiance function of the 3D Gaussians to address these limitations.
@@ 3.2 Disentanglement of Isotropic and Anisotropic 3D Gaussians
We argue that, while X-ray scattering models are incompatible with the 3DGS rendering formulation (secondary rays would have to be cast, increasing the computational footprint exponentially), their anisotropic impact on X-ray imaging can be approximated to some extent by direction-dependent radiance functions. However, modeling this high-dimensional residual contribution is only meaningful if it appears in the training DRRs, i.e., if the target radiographs are rendered using physics-inspired simulation tools. This is not always possible, e.g., if the mapping from HU values to materials for the target CT scans is not provided. Ideally, in such cases, when only scatter-free DRRs can be generated, the radiance function should gracefully degrade to a lighter anisotropic formulation (e.g., similar to Cai et al. (2024)), to avoid unnecessary computations that could impact the model optimization, as well as its runtime performance. Moreover, X-ray interactions with matter themselves involve both isotropic (e.g., photoelectric absorption/fluorescence, Rayleigh scattering) and anisotropic (e.g., Compton scattering) phenomena Hasegawa (1990); Migliori . Therefore, for the sake of both modularity and accuracy, we propose to decompose our DRR representation into isotropic Gaussians $g_{i}^{\text{iso}}$ and anisotropic, direction-dependent ones $g_{j}^{\text{dir}}$. We formulate their respective radiosity functions as:
@@
where $\bm{f}_{i}^{\text{iso}},\bm{f}_{j}^{\text{dir}}\in\mathbb{R}^{k}$ are feature vectors respectively encoding the contribution of isotropic and anisotropic Gaussians, $\bm{b}^{\text{iso}}\in\mathbb{R}^{k}$ is a global isotropic basis vector (similar to Cai et al. (2024)), and $\bm{B}^{\text{dir}}\in\mathbb{R}^{k_{L}\times k}$ is a second, anisotropic basis matrix that interacts with the $L$-degree spherical-decomposition $Y:\mathbb{R}^{2}\mapsto\mathbb{R}^{k_{L}}$ applied to the ray angles $\theta,\phi$ without constant term (degree $=$ 0), resulting in $k_{L}=L(L+2)$.
@@ 3.3 Initialization via Radiodensity-Aware Dual Sampling (RADS)
Our second main contribution targets the initialization of the point-cloud that supports the Gaussian-mixture modeling of the 3D data. Since 3DGS was defined for natural imaging (e.g., ignoring sub-surface light transport), common initialization strategies limit their sampling to scene-surface points (SfM- Kerbl et al. (2023) or depth-guided Niemeyer et al. (2024) subsampling). Such sampling of the 3D space is, however, inadequate for volumetric data, ignoring possibly salient regions. While uniform sampling of the CT space could be considered Cai et al. (2024); Nikolakakis et al. (2024), we argue that such a strategy is also suboptimal for anatomical data, discarding domain-relevant properties. We demonstrate empirically (c.f. Section 4) that uniform sampling results in slower convergence of the 3DGS representation and degraded runtime performance.
@@
Not unlike Niemeyer et al. (2024), the first set of points is sampled by applying the marching-cubes algorithm Lorensen and Cline (1998) to extract points at the interface between materials with different physical properties (i.e., distinct HU responses in the CT volume). These interface regions typically need careful modeling when simulating X-ray imaging, as they result in discontinuous photon transport.

### v2

@@ 1 Introduction
Digitally reconstructed radiographs (DRRs) are simulated (in silico) 2D X-ray images rendered from 3D computational tomography (CT) volumes. While DRRs are widely utilized in preoperative settings, such as optimizing dose delivery and in radiation oncology [20, 41, 6], their potential intraoperative applications remain underexplored due to computational bottlenecks and naive modeling capability [1, 27]. For instance, real-time multimodal registration for image-guided procedures is currently impractical because of the time-consuming process of generating DRRs and integrating them into slice-to-volume registration [36].
The advent of GPU-accelerated computing has significantly improved the efficiency of radiography simulators [3, 39, 5], though they still fall short of meeting the stringent requirements of real-time applications. Recent developments in inverse graphics have led to the creation of differentiable DRR renderers (e.g., DiffDRR [11], X-Gaussian [7], GaSpCT [25]), which show promise for inverse problems such as 2D/3D CT image registration [12]. However, these methods do not fully capture the complexities of X-ray image formation. They typically use computationally efficient ray-tracing techniques that model the attenuated photon fluence at each detector pixel by accumulating Hounsfield unit (HU) values along the 3D view-ray through the voxelized CT volume. This approach, while fast, cannot account for complex noise-inducing physical effects such as Compton scattering and beam hardening [14], crucial for accurate X-ray imaging.
Physics-based Monte-Carlo simulations [2, 15, 3, 4] do not suffer from these limitations. Relying on the decomposition of CT volumes into realistic materials via HU thresholding, they accurately simulate single-photon transport and probabilistically evaluate photon-matter interactions (photoelectric effect, Compton scattering, Rayleigh scattering, etc.) to determine the final attenuation and hit-region of each photon. This approach provides a more accurate representation of X-ray image formation; but the computational intensity of these methods—requiring >$10^{8}$ iterations to render an image—makes them impractical for real-time applications and inverse problems.
@@
Our approach leverages 3D Gaussian splatting (3DGS) [17] for efficient DRR generation from CT volumes, similar to X-Gaussian [7] and GaSpCT [25]. However, we introduce a crucial adaptation: our direction-disentangled 3DGS (DDGS) method accounts for the dependencies of X-ray image formation on light directionality. Specifically, we decompose the radiosity contribution of 3D regions into two independent components: a non-directional component modeling the isotropic X-ray interactions with the corresponding material, and a direction-dependent component approximating higher-dimensional anisotropic interactions, such as scattering. Unlike the more generic view-dependent spherical-harmonics decomposition adopted by common 3DGS solutions [10], ours is tuned to the specific behaviors of high-energy photons. This formulation allows our method to implicitly learn and reproduce the effects of scattering on X-ray images without the need for complex photon-transport simulations at runtime.
@@ 2 Related Work
The generation of DRRs and the simulation of X-ray images have been subjects of extensive research due to their critical roles in medical imaging, particularly in radiotherapy planning and diagnostic imaging [20, 41, 6, 43]. In this section, we provide some insight on X-ray imaging, traditional simulation tools, and the recent integration of machine-learning techniques.
@@
X-ray imaging relies on the interactions of high-energy photons—emitted by the X-ray source of the scanner—with the matter (e.g., body part) placed in between the X-ray source of the scanner and its radiation-sensitive detector cells (dexels). Imaging contrast arises from differential absorption and scattering of X-ray photons by tissues, primarily through the photoelectric effect, Compton scattering, and, to a lesser extent, beam hardening for polychromatic scanners. The photoelectric effect involves the absorption of an X-ray photon by an atom and is more pronounced in higher atomic number materials like bone, causing them to appear whiter on X-ray images [^1] X-ray film is historically white, turning black when exposed to high-energy photon.; whereas Compton scattering occurs when an X-ray photon collides with an outer electron, resulting in the photon being deflected and losing energy, which contributes to image noise and reduced contrast [14, 29].
A variety of Monte Carlo (MC) simulation suites [2, 15, 3, 4] model these X-ray interactions and light transport. They trace the paths of a large number of individual photons as they undergo probabilistic interactions within the material. By simulating numerous photon trajectories, Monte Carlo methods account for the random nature of photoelectric absorption and Compton scattering, as well as secondary effects such as fluorescence and electron transport. These MC solutions produce highly realistic simulations of radiographic images, though at a high computational cost which makes them inadequate for most DRR use-cases. Furthermore, for these tools to render realistic DRRs from CT volumes, one must first pre-process the CT data to assign proper materials and their corresponding physical properties to each voxel, which is essential for realistic simulation. This step is often approximated by decomposing the volume into a small set of predefined materials according to user-provided HU ranges (e.g., “air” material assigned to voxels with HU value $<-800$, “lung” assigned to voxels in range ]800,200]]-800,-200]] - 800 , - 200 ], “fat” for ]200,100]]-200,-100]] - 200 , - 100 ], etc.). This process can be time-consuming and prone to errors, affecting the fidelity of MC-based DRR generation [9].
@@
Analytical DRR rendering methods [39, 38, 5] have been developed in parallel to the aforementioned MC suites for real-time or near-real-time applications. These methods have a much lower computational footprint, as they do not simulate individual photon interactions but instead use mathematical models to directly calculate the primary paths and approximated attenuations of X-rays. The seminal work of Siddon [31] on computing line integrals through a discretized CT volume is still at the core of most recent DRR solutions, e.g., which proposed differentiable formulation of the rendering steps [11] or integrated neural networks to add realistic noise to the output [35]. However, the inherent approximations in these methods can affect their accuracy, particularly in modeling artifacts present in heterogeneous tissues or complex anatomical structures, e.g., caused by scattering (non-primary light contributions).
Recently, researchers [7, 25] have tried to apply 3D Gaussian splatting (3DGS) techniques [17] to DRR rendering, i.e., optimizing a cloud of 3D Gaussians to approximate voxel data, thereby enabling faster rendering with minimal accuracy loss. However, these early solutions overlooked the specific nuances of X-ray imaging, directly applying 3DGS methods designed for natural imaging [25] or oversimplifying the physical properties of X-ray attenuation [25], leading to suboptimal modeling.
@@ 3.1 Preliminaries
3DGS [17] is a point-based rendering technique using 3D Gaussians to explicitly represent scenes in a more compact manner than volumetric representations. Each Gaussian $g_{i}=\{\bm{\mu}_{i},\bm{\Sigma}_{i},\bm{\alpha}_{i},\bm{f}_{i}\}$ is defined by its mean position $\bm{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ (usually decomposed into separate scale and rotation parameters), opacity $\bm{\alpha}_{i}\in\mathbb{R}$, and radiance properties $\bm{f}_{i}\in\mathbb{R}^{k}$, where $k$ depends on the light contribution model adopted. E.g., $k=(L+1)^{2}\times 3$ for anisotropic, view-dependent RGB radiance decomposed into spherical harmonics (SH) of degree $L$, a model commonly used for natural-imaging applications. Images are rendered by casting view-rays through each pixel $p$ into the scene and alpha-blending the Gaussian contributions to the final ray color $C$, as:
@@
where $n$ is the number of Gaussians considered, $c_{j}$ is the radiance (computed from $\bm{f}_{j}$, e.g., via SH) of the $j$th Gaussian on the ray, and $\widehat{\bm{\mu}}_{j}$ and $\widehat{\bm{\Sigma}}_{j}$ are the image-plane projections of $\bm{\mu}_{j}$ and $\bm{\Sigma}_{j}$.
Since the aforementioned rasterization process is differentiable w.r.t. to the Gaussian parameters, the 3DGS representation of a scene can be learned via gradient-descent, given a set of 2D scene observations $\bm{I}_{\text{GT}}$ and their corresponding camera parameters. A combination of $\ell_{1}$ and SSIM [40] loss functions are commonly-adopted as criterion for this iterative optimization process [17]:
@@
with $\bm{I}$ the rendered images (based on provided camera parameters), and $\lambda$ a loss-weighting hyper-parameter. The initial state is typically obtained by sampling relevant 3D points in the scene domain for the Gaussians, e.g., from a structured-from-motion (SfM) point-cloud [34, 28].
@@
The volumetric rendering performed by 3DGS methods is conceptually similar to the one performed by analytical DRR methods [31, 11], i.e., aggregating the attenuation values of light-rays cast through the voxelized CT volume. This similarity has recently motivated researchers [7, 25] to apply 3DGS models to DRR applications, driven by the compactness of 3DGS representations compared to voxel data (addressing the $\mathcal{O}(n^{3})$ complexity inherent to voxel rendering). By optimizing a 3DGS model to approximate the visual properties of the CT volume, the resulting representation can render accurate DRRs much faster than traditional methods, making it more suitable for real-time intra-operative visualization. Consequently, the use of traditional, slower rendering solutions can be limited to generating DRR targets ($\bm{I}_{\text{GT}}$) for the offline (i.e., pre-operative) training.
However, these early 3DGS-for-DRR solutions [7, 25] do not account for noise-inducing photon interactions (e.g., scattering) when applying the analytical methods [30, 5] to render their training data. Since scatter-free DRRs are not affected by ray directions (i.e., the attenuation at each point is naively considered isotropic), Cai et al. [7] simplified the radio-intensity function of each Gaussian to $c_{i}=\operatorname{sigmoid}{(\bm{b}\cdot\bm{f}_{i})}$, where $\bm{b}\in\mathbb{R}^{k}$ is a direction-independent optimizable basis vector shared by all splats. Furthermore, their Gaussian point-cloud is initialized via evenly-spaced Cai et al. [7] or random-uniform [25] sampling of the scanned 3D space, i.e., ignoring relevant geometrical and material properties of the target CT data. Inadequate initialization strategies negatively impact the convergence of 3DGS models and the resulting image accuracy [16]. In this work, we propose to reformulate both the initialization strategy and the radiance function of the 3D Gaussians to address these limitations.
@@ 3.2 Disentanglement of Isotropic and Anisotropic 3D Gaussians
We argue that, while X-ray scattering models are incompatible with the 3DGS rendering formulation (secondary rays would have to be cast, increasing the computational footprint exponentially), their anisotropic impact on X-ray imaging can be approximated to some extent by direction-dependent radiance functions. However, modeling this high-dimensional residual contribution is only meaningful if it appears in the training DRRs, i.e., if the target radiographs are rendered using physics-inspired simulation tools. This is not always possible, e.g., if the mapping from HU values to materials for the target CT scans is not provided. Ideally, in such cases, when only scatter-free DRRs can be generated, the radiance function should gracefully degrade to a lighter anisotropic formulation (e.g., similar to [7]), to avoid unnecessary computations that could impact the model optimization, as well as its runtime performance. Moreover, X-ray interactions with matter themselves involve both isotropic (e.g., photoelectric absorption/fluorescence, Rayleigh scattering) and anisotropic (e.g., Compton scattering) phenomena [14, 22]. Therefore, for the sake of both modularity and accuracy, we propose to decompose our DRR representation into isotropic Gaussians $g_{i}^{\text{iso}}$ and anisotropic, direction-dependent ones $g_{j}^{\text{dir}}$. We formulate their respective radiosity functions as:
@@
where $\bm{f}_{i}^{\text{iso}},\bm{f}_{j}^{\text{dir}}\in\mathbb{R}^{k}$ are feature vectors respectively encoding the contribution of isotropic and anisotropic Gaussians, $\bm{b}^{\text{iso}}\in\mathbb{R}^{k}$ is a global isotropic basis vector (similar to [7]), and $\bm{B}^{\text{dir}}\in\mathbb{R}^{k_{L}\times k}$ is a second, anisotropic basis matrix that interacts with the $L$-degree spherical-decomposition $Y:\mathbb{R}^{2}\mapsto\mathbb{R}^{k_{L}}$ applied to the ray angles $\theta,\phi$ without constant term (degree $=$ 0), resulting in $k_{L}=L(L+2)$.
@@ 3.3 Initialization via Radiodensity-Aware Dual Sampling (RADS)
Our second main contribution targets the initialization of the point-cloud that supports the Gaussian-mixture modeling of the 3D data. Since 3DGS was defined for natural imaging (e.g., ignoring sub-surface light transport), common initialization strategies limit their sampling to scene-surface points (SfM- [17] or depth-guided [24] subsampling). Such sampling of the 3D space is, however, inadequate for volumetric data, ignoring possibly salient regions. While uniform sampling of the CT space could be considered [7, 25], we argue that such a strategy is also suboptimal for anatomical data, discarding domain-relevant properties. We demonstrate empirically (c.f. Section 4) that uniform sampling results in slower convergence of the 3DGS representation and degraded runtime performance.
@@
Not unlike [24], the first set of points is sampled by applying the marching-cubes algorithm [21] to extract points at the interface between materials with different physical properties (i.e., distinct HU responses in the CT volume). These interface regions typically need careful modeling when simulating X-ray imaging, as they result in discontinuous photon transport.